2. Calculate the molar solubility of Ag:CrOs at 25 °C given that the Kip for the salt is 2.0x10 3. Show if a precipitate of PbCl2 will form if you mix 5.0 ml of 0.001 M Pb2+ion solution with 5.0 mL of 0.0015 M solution of CI. Ksp of PbCl2 1.7x 10
Answer #1

Ans 2

The dissociation reaction

Ag2CrO4 = 2Ag+ + CrO4(-2)

Ksp = [Ag+]^2 [CrO42-]

let x = mol/L

Molarity of Ag2CrO4 = x

Molarity of Ag+ = 2x

Molarity of CrO42- = x

2 * 10^-12 = (2x)^2 * x

4x^3 = 2*10^-12

x = 8.722 * 10^-5 = molar solubility of Ag2CrO4